A research was conducted in 2017 at the Tema Port to find the ages of cars imported into 

the country. According to the research, 10% of the cars imported were less than one-year￾old. Assuming this result holds true for the current period for all cars imported into the 

country.

Find the probability that in a random sample of 5 cars at the Tema Port

(i) exactly 3 are less than one-year-old.

(ii) exactly 2 are less than one-year-old.

(iii) none is less than one-year-old 

(iv) Find the mean and standard deviation of the distrib


1
Expert's answer
2021-04-15T06:57:20-0400

This is a binomial distribution with p=0.1, n=5.


(i) "P(X=3)=C_5^30.1^3(1-0.1)^2=0.0081."


(ii) "P(X=2)=C_5^20.1^2(1-0.1)^3=0.0729."

(iii) "P(X=0)=(1-0.1)^5=0.5905."


(iv) "\mu=np=5*0.1=0.5."

"\sigma=\sqrt{np(1-p)}=\sqrt{0.45}=0.6708."